Problem 3-4A (Video) (Part Level Submission)

Rivera Company has several processing departments. Costs charged to the Assembly Department for November 2020 totaled $2,313,896 as follows.

Work in process, November 1
   Materials $78,800
   Conversion costs 48,100 $126,900
Materials added 1,622,970
Labor 225,100
Overhead 338,926


Production records show that 34,900 units were in beginning work in process 30% complete as to conversion costs, 659,700 units were started into production, and 24,100 units were in ending work in process 40% complete as to conversion costs. Materials are entered at the beginning of each process.

Determine the assignment of costs to goods transferred out and in process.

Costs accounted for:

   Transferred out

$

   Work in process, November 30

      Materials

$

      Conversion costs

         Total costs

$

2) create a production cost report

3) Cost Reconciliation Schedule
Answer #1
1
Costs accounted for:
Transferred out 2246175
Work in process, November 30
      Materials 59045
      Conversion costs 8676 67721
Total costs 2313896
2
RIVERA COMPANY
Assembly Department
Production Cost Report
For the Month Ended November 30, 2020
Equivalent Units
Physical Units Materials Conversion Costs
Units to be accounted for
Work in process, November 1 34900
Started into production 659700
Total units 694600
Units accounted for
Transferred out 670500 670500 670500
Work in process, November 30 24100 24100 9640
Total units 694600 694600 680140
Materials Conversion Costs Total
Unit costs:
Total costs 1701770 612126 2313896
Equivalent units 694600 680140
Unit costs 2.45 0.90 3.35
Costs to be accounted for
Work in process, November 1 126900
Started into production 2186996
Total costs 2313896
Cost Reconciliation Schedule
Costs accounted for
Transferred out 2246175 =670500*3.35
Work in process, November 30
      Materials 59045 =24100*2.45
      Conversion costs 8676 67721 =9640*0.90
Total costs 2313896